Displaying report 1-1 of 1.
Reports until 16:59, Friday 19 July 2024
GQuEST Output Filter Cavities
Torrey Cullen - posted 16:59, Friday 19 July 2024 (11761)
FC2/3 sled update

FC2/3 Sled is almost prepped for super optics. 1550 light is ready to be mode matched. The twice shifted via AOM 775 light is successfully fed to the FC2/3 sled via fiber. There is a 5 meter fiber for both wavelengths. On the sled a 2 m fiber for the 775 path and a 50:50 fiber BS for the 1550 path, as seen in this. The same will be done for 775 light for the third filter cavity. Tasks to be done:

-Mode match both wavelengths.

-Install PDs, wire up, etc.

-Assemble cavity optomechanics (flexture mounts)

-Superoptics.

Images attached to this report
Displaying report 1-1 of 1.